# Testing Dark Matter with Generative Models for Extragalactic Stellar Streams

### v1

@@ 6 Summary & Conclusion
Upcoming and ongoing surveys including Euclid (Racca et al., 2016), The Rubin Observatory (Ivezic et al., 2008; Ivezić et al., 2019), Roman (Spergel et al., 2015), and ARRAKIHS (Guzman et al., 2022) are expected to discover 1000s of extragalactic stellar streams. The method presented here provides a generative approach to transform stream images into constraints on the underlying mass profile of dark matter halos. Because streams often reside in the outer regions of dark matter halos, the method provides independent constraints on halo shapes and mass limits, and probes a region of the halo that is inaccessible to traditional techniques using stellar and HI kinematics. The method is applicable at the level of individual galaxies, while most previous applications have been limited to galaxy groups or clusters, utilizing X-ray emission (e.g. Reiprich et al., 2013) or weak-lensing (e.g. Umetsu et al., 2011). Similar to lensing, it is possible to combine information across many streams or hosts to obtain a higher confidence constraint on halo profiles. Combined with recent work showing that the shapes and masses of dark matter halos can be inferred from streams (Fardal et al., 2013; Pearson et al., 2022b; Nibauer et al., 2023; Walder et al., 2024), and that gaps from dark matter subhalos can be observed in extragalactic globular cluster tidal tails (Aganze et al., 2024), extragalactic stream imaging will offer a powerful test of dark matter across 1000s of galaxies spanning a wide range of masses and redshifts.
@@ Appendix B Validation of Test Statistic
Figure 11: Validation that the derived $\bar{\Lambda}$ from Eq. 9 (dashed red curve) matches $\Lambda$ values measured from many simulations (solid black curve). Under the assumptions of normality, one can map $\Lambda$ to a $\Delta\sigma$ deviation by numerically inverting Eq. 9.
Here we validate the test statistic, $\Lambda$, derived in §3.4.1. Under the assumptions of normality discussed in §3.4.1, we generate samples from two normals and measure the test statistic $\Lambda$ given by Eq. 6. We perform simulations as a function of the parameter $\sigma=\Delta\sigma+1$. We than compute analytically Eq. 9, and compare the predicted values of $\Lambda$ to the derived values from the simulations in Fig. 11. Indeed, there is excellent agreement between the simulated (black) and predicted (dashed red) curves, indicating that our test statistic can be used to map from $\Lambda\xrightarrow[]{}\Delta\sigma$.
@@ Appendix C Outer Stream Constraint
[…]

### v2

@@ VI Summary & Conclusion
Upcoming and ongoing surveys including Euclid (Racca et al., 2016), The Rubin Observatory (Ivezic et al., 2008; Ivezić et al., 2019), Roman (Spergel et al., 2015), and ARRAKIHS (Guzman et al., 2022) are expected to discover 1000s of extragalactic stellar streams. The method presented here provides a generative approach to transform stream images into constraints on the underlying mass profile of dark matter halos. Because streams often reside in the outer regions of dark matter halos, the method provides independent constraints on halo shapes and mass limits, and probes a region of the halo that is inaccessible to traditional techniques using stellar and HI kinematics. The method is applicable at the level of individual galaxies, while most previous applications have been limited to galaxy groups or clusters, utilizing X-ray emission (e.g. Reiprich et al., 2013) or weak-lensing (e.g. Umetsu et al., 2011). Similar to lensing, it is possible to combine information across many streams or hosts to obtain a higher confidence constraint on halo profiles. Combined with recent work showing that the shapes and masses of dark matter halos can be inferred from streams (Fardal et al., 2013; Pearson et al., 2022b; Nibauer et al., 2023; Walder et al., 2025), and that gaps from dark matter subhalos can be observed in extragalactic globular cluster tidal tails (Aganze et al., 2024), extragalactic stream imaging will offer a powerful test of dark matter across 1000s of galaxies spanning a wide range of masses and redshifts.
@@ Appendix B Validation of Test Statistic
Figure 11: Validation that the derived $\bar{\Lambda}$ from Eq. 9 (dashed red curve) matches $\Lambda$ values measured from many simulations (solid black curve) under the assumption of normality. We also compute $\Lambda\left(\Delta\sigma\right)$ for a Student’s t-distribution with $\nu$ degrees of freedom. At small $\nu$ (heavier tails) the true $\Lambda$ values are higher than the Gaussian prediction.
Here we validate the test statistic, $\Lambda$, derived in §III.4.1. Under the assumptions of normality discussed in §III.4.1, we generate samples from two normals and measure the test statistic $\Lambda$ given by Eq. 6. We perform simulations as a function of the parameter $\sigma=\Delta\sigma+1$. We than compute analytically Eq. 9, and compare the predicted values of $\Lambda$ to the derived values from the simulations in Fig. 11. Indeed, there is excellent agreement between the simulated (black) and predicted (dashed red) curves under the assumption of normality. To illustrate departures from a Gaussian distribution, we utilize a Student’s t-distribution with $\nu$ degrees of freedom. We consider $\nu\in\{4,7,20\}$, noting that as $\nu\xrightarrow{}\infty$, the distribution approaches a standard normal. For each choice of $\nu$, we generate samples and adjust the scale parameter, $s$, such that $x\xrightarrow{}x/s$. For every combination of $\nu$ and $s$, we compute the area under the curve from the center of the distribution up to the 34th percentile. This allows us to derive the $\Delta\sigma$ factor relative to the baseline case where $s=1$. Finally, we compute the statistic $\Lambda$, with the resulting values illustrated in Fig. 11 (blue, orange, and green curves for $\nu=4,7,$ and $20$, respectively). At $\nu=4$, the Student’s t-distribution is heavy-tailed compared to the normal distribution, resulting in a $\Lambda$ that is larger than the prediction under normality (dashed red curve). As $\nu$ increases, this discrepancy lessens. The discrepancy between the true $\Lambda(\Delta\sigma)$ curves and the normal prediction is maximal at large $\Delta\sigma$. However, in this work, we draw contours using $\Lambda(\Delta\sigma<2)$. Fig. 11 illustrates that at this level, our contours can be viewed as conservative, since heavy-tailed distributions produce more positive $\Lambda$ values than a Gaussian distribution at a fixed $\Delta\sigma$.
@@ Appendix C Outer Stream Constraint
In Fig. 12 we present the results of our sampler for the outer stream. The purple contours show the 68% and 95% confidence regions. The true parameters are over plotted as black lines. The colored points represent solutions which fall within the $68\%$ region (blue and red), and a discrepancy solution (green) (see Fig. 6 for the corresponding stream models). All 10 free parameters for the progenitors and the dark matter halos are recovered within the $68\%$ region, though the stream morphology is not sensitive to the scale radius, $r_{s}$, where the posterior distributions are flat, and we can only place a lower limit on the dynamical age of the stream, $t_{\rm age}$. Compared to the inner stream (see Fig. 5), the outer stream places a weaker constraint on the inner density slope, $\gamma$, but a stronger constraint on the outer density slope, $\beta$ (see the discussion of Fig. 7 in §IV).
@@
Appendix D Cored Stream Constraint
In Figure 9, the sampler recovers the preference for a cored profile and the true parameters are within 68% contours. To explore valid fits, we select three different model streams within the 68% contours from Figure 9 in the 10 dimensional space, which have $\beta$ = 2.2, 2.25, 2.3 and $0.15<\gamma<0.3$. Note that the other 8 parameters are selected at random within 68% contours and vary in each case. We visualize these three model streams in Figure 13. While all streams produce good fits to the input data, there are some differences between each stream. The left stream with the lowest $\gamma$ and $\beta$ is slightly longer and wider than the other streams. This stream also happens to be from a more massive progenitor fit within the 68% contours compared to the middle and right stream.
Figure 13: Three different streams (black points) which are evolved in a halo with low $\gamma$ values (see legends) selected within the 68% contours in 10 dimensions from Figure 9. The red points show the input data used to represent the inner stream, and all streams produce good fits to the input data.